What is the equation in point-slope form for the line parallel to y = –6x + 1 that contains J(–6, 8)?
Answer by mananth(16949)

6 x + 1 y = 1
Find the slope of this line
make y the subject
1 y = -6 x + 1
Divide by 1
y = -6 x + 1
Compare this equation with y=mx+b
slope m = -6
The slope of a line parallel to the above line will be the same
The slope of the required line will be -6
m= -6 ,point ( -6 , 8 )
Find b by plugging the values of m & the point in
y=mx+b
8 = 36 + b
b= -28
m= -6
Plug value of the slope and b in y = mx +b
The required equation is y = -6 x -28
m.ananth@hotmail.ca